### v2

@@ Abstract
In this paper, we study a class of real-valued mean-field backward stochastic differential equations (BSDEs) with generator of quadratic growth in the control variable and the mean-field term. Under this assumption, together with a bounded terminal condition, we establish existence and uniqueness of solutions. Our approach departs from classical fixed-point arguments and instead combines Malliavin calculus with refined BMO and stability estimates. The result bridges the gap between the quadratic BSDE results of cheridito2017bse and hao2022mean. Moreover, motivated by the structure of the mean-field exponential utility maximization problem introduced in our paper, we extend our framework to generators satisfying a weaker quadratic condition on the generator. This relaxation is designed to accommodate the additional mean-field terms that arise in our utility maximization setting and that fall outside the scope of previous quadratic assumptions. Within this more general regime, we establish existence and uniqueness of solutions under a smallness condition on the terminal random variable. We then apply this extended theory to solve a mean-field exponential utility maximization problem, thereby generalizing the classical framework of hu2005utility to a fully coupled quadratic mean-field setting.
Key words: mean-field BSDEs; quadratic BSDEs; BMO martingales; Malliavin calculus; exponential utility maximization.
@@ 1 Introduction
The origin of linear BSDEs can be traced back to their role as adjoint equations in stochastic control problems, first studied by bismut1973conjugate. The seminal paper by pardoux1990adapted established the existence and uniqueness of adapted solutions $(Y,Z)$ for
@@
under the conditions that the generator $F$ is Lipschitz continuous and the terminal condition $\xi$ is square integrable. Since then, there has been a growing interest in BSDEs with applications in various fields such as stochastic control, mathematical finance, and partial differential equations. For a comprehensive presentation of the theory, see el1997backward, CohenStocal2015 or ZhangBSDEbook2017. Numerous studies have aimed to improve existence and uniqueness results by relaxing the Lipschitz condition on the generator $F$ or the conditions on the terminal condition $\xi$.
@@
BSDEs with generators exhibiting quadratic growth in the variable $z$ have attracted significant attention. These equations arise, for example, in stochastic linear-quadratic control with random coefficients bismut1976linearquad, in utility optimization problems in an incomplete market hu2005utility. Kobylanski:2000cy provided the seminal existence and uniqueness result for real-valued quadratic BSDEs with bounded terminal values, where $F$ has quadratic growth with respect to its $Z$ component. briand2013simple have continued to explore this area, demonstrating the existence of unique solutions to quadratic BSDEs with bounded terminal conditions using purely probabilistic methods. Investigations by briand2006bsde extend the previous results to unbounded terminal conditions. Some studies the multidimensional cases: see hu2016multi; tang2023MultiMFBSDE; luo2019triangle; tevzadze2008solvability.
@@
Mean-field stochastic differential equations, also known as McKean-Vlasov equations, describe stochastic systems whose evolution is influenced by both the microscopic state and the macroscopic distribution of particles. These equations date back to kac1956foundations’s work on the Boltzmann and Vlasov equations in the 1950s. Mean-field game theory began with the groundbreaking contributions of lasry2007mfgame. Building on this foundation, Buckdahn2009MFlimit introduced mean-field BSDEs, derived from approximations involving weakly interacting $N$ particle systems, and investigated them using probabilistic methods. This has led to a growing interest in mean-field BSDEs and FBSDEs. For instance, Buckdahn2009MFBSDE studied the existence and uniqueness of solutions with Lipschitz continuous generators and square integrable terminal conditions. More recently, hao2022mean explored the solvability of solutions for one-dimensional mean-field BSDEs with quadratic growth, while tang2023MultiMFBSDE established the existence and uniqueness of solutions for multidimensional BSDEs with generators depending on the expectation of both variables. Li2018MFBSDEcontinuous examined mean-field BSDEs whose coefficients depend not only on the solution $(Y,Z)$ but also on the law of $Y$.
@@
pardoux1990adapted interpreted the key structural feature of BSDEs as a nonlinear conditional expectation, highlighting their close link to Malliavin calculus. In fact, el1997backward and Buckdahn2009MFBSDE state that, in a Clark–Ocone type formula, the martingale density component $Z$ of the solution pair of a BSDE with a classical globally Lipschitz generator on a Gaussian basis is the Malliavin derivative of the other component $Y$ (see Proposition 5.3). The result has been extending our understanding of quadratic BSDEs: for example, briand2013simple provided a straightforward approach to the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus to overcome difficulties with quadratic generators, and Imkeller2010path proposed a numerical scheme for BSDEs with generators of quadratic growth.
@@
Continuous time portfolio optimization under the utility-maximization framework introduced by merton1969portfolio is a foundational topic in mathematical finance. Merton analyzed the problem in a complete market where the risky asset follows the Black-Scholes dynamics and no constraints are imposed on admissible trading strategies. A major development in the early 2000s was the emergence of a connection between this class of optimization problems and the theory of BSDEs, which has since become a central approach in the field. The core idea of the BSDE methodology is to construct a stochastic process, dependent on the investment strategy, that coincides with the investor’s utility of terminal wealth. In this context, elrouge2000utility used BSDEs to compute the value function and optimal strategy for exponential utility, assuming trading strategies are restricted to a convex cone. Building on this framework, hu2005utility developed a BSDE approach to utility maximization in incomplete markets and extended the methodology to include logarithmic and power utility functions, allowing for general strategy constraints within a closed set.
@@
Our methods are inspired by the works of tevzadze2008solvability and briand2013simple. The former studied the existence and uniqueness of solutions for multidimensional quadratic BSDEs, utilizing BMO martingale theory. The latter provided a straightforward approach to establishing the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus. Building on these foundations, we bridge the gap between the approaches of cheridito2017bse and hao2022mean. Specifically, hao2022mean established the existence and uniqueness by applying a contraction mapping with the Reverse Hölder Inequality (RHI) and iterating backward in time when the generator is quadratic in $Z$ component, while cheridito2017bse used the Krasnoselskii fixed-point theorem on a small time interval, iterating backward to obtain a global solution. Our contribution extends the results of cheridito2017bse by demonstrating in Proposition 4.9 that the function $f^{1}$ satisfies the conditions outlined in hao2022mean, thereby proving the existence and uniqueness of a solution for quadratic mean-field BSDEs, which is quadratic both in $Z$ and the corresponding mean-field component $\tilde{Z}$.
Our main contributions are as follows. First, we establish the existence and uniqueness of a solution when the generator exhibits separately quadratic growth in both $Z$ and the mean-field term $\tilde{Z}$ with bounded terminal condition $\xi$. Our approach builds upon the theory of BMO martingales and Malliavin calculus to effectively address the analytical challenges arising from the quadratic structure. A key novelty of this work lies in the inclusion of a quadratic mean-field component $\tilde{Z}$, which distinguishes it from existing results in the literature. Moreover, our method is notable for its simplicity, as it avoids the use of fixed-point arguments commonly employed in previous studies, such as juan2024MFBSDE and tevzadze2008solvability.
@@
The rest of the paper is organized as follows. In Section 2, we present the mathematical framework and state the main result. Section 3 derives an almost sure upper bound for the solution pair $(Y,Z)$ of Lipschitz mean-field BSDEs, assuming the terminal condition admits a uniformly bounded Malliavin derivative, as shown in Proposition 3.1. In the main Section 4, we provide a priori estimates and establish a powerful stability result for mean-field BSDEs with separately quadratic growth. In Section 5, we utilize the bounds obtained in Section 3, which are independent of the Lipschitz constants in $z$ and $\tilde{z}$, together with the stability result and a density argument, to prove the existence and uniqueness of solutions to our quadratic mean-field BSDEs. Finally, in Section 6, we extend the existence and uniqueness results for mean-field quadratic BSDE with separately quadratic driver to a fully coupled quadratic driver with small terminal conditions, and show how the corresponding BSDE characterizes a utility maximization problem with a particular class of constraint sets.
@@ 2.1 Preliminaries
  • $\mathbb{H}^{p}$: the set of predictable $\mathbb{R}^{d}$-valued process $Z$ satisfying $\left\|Z\right\|_{\mathbb{H}^{p}}:=\left(\mathbb{E}\left[\int_{0}^{T}\left|Z_{r}\right|^{2}\,dr\right]^{\frac{p}{2}}\right)^{\frac{1}{p}}<\infty,$
@@
    $\mathbb{E}\left[\int_{\tau}^{T}\left|Z_{r}\right|^{2}\,dr\Big|\mathcal{F}_{\tau}\right]\leq C^{2}.$
    The $\left\|Z\right\|_{\mathbb{H}^{2}_{BMO}}:=\left\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\right\|_{BMO}$ is defined as the smallest non-negative constant $C$ for which the above inequality is satisfied.
@@
When a different underlying measure other than $\mathbb{P}$ is involved, we indicate it explicitly in the notation, for example by writing $\mathbb{E}$$^{\mathbb{Q}}$$\cdot$] and $\|\cdot\|_{BMO(\mathbb{Q})}$ for expectations and BMO norms taken with respect to a probability measure $\mathbb{Q}$. Here are some inequalities from BMO martingales theory, all the proofs can be found in kazamaki2006BMO.
@@
For $K>0$. There exist constants $c_{1}(K),c_{2}(K)>0$, depending only on $K$, such that the following holds. Let $N$ be a one-dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$ with $\|N\|_{\mathrm{BMO}(\mathbb{P})}\leq K$, and define the equivalent measure $\tilde{\mathbb{P}}$ by
@@
For any continuous $\mathrm{BMO}$ martingale $M$ under $\mathbb{P}$, set
$\tilde{M}\;:=\;M-\langle M,N\rangle,$
@@ 1. $(\rm H_{0})$}
  2. $(\rm H_{q})$}
@@
The following $g$ satisfies the condition $(\mathrm{H}_{q})$:
@@
where $l$ is Lipschitz in (z,z~)($\tilde{z}$), and $q_{1}$ and $q_{2}$ are quadratic polynomials in $z$ and $\tilde{z}$, respectively. Note that this example does not satisfy the existence and uniqueness assumption in previous literature, such as hao2022mean.
@@
Let Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$ hold. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
@@
In the statement and proof of our main theorem we may, without loss of generality, we can assume that
@@
This is not an additional structural condition on the original generator, but merely a convenient normalisation. Given any generator $g$ satisfying Assumptions $(\rm H_{0})$$(\rm H_{q})$}, we may introduce the shifted generator
@@
$Y_{t}:=U_{t}+\int_{t}^{T}\widetilde{{\mathbb{E}}}\big[g(s,0,\tilde{Z}_{s})\big]\,ds$

### v3

@@ Abstract
In this paper, we study a class of real-valued mean-field backward stochastic differential equations (BSDEs) with generators of quadratic growth in the control variable and the mean-field term. Under this assumption, together with a bounded terminal condition, we establish the existence and uniqueness of solutions. Our approach departs from classical fixed-point arguments and instead combines Malliavin calculus with refined BMO and stability estimates. The result bridges the gap between the quadratic BSDE results of [Ann. Probab. 45 (2017), pp. 3795–3828] and Hao et al. [Ann. Appl. Probab. 35 (2025), pp. 2128–2174]. Moreover, motivated by the structure of the mean-field exponential utility maximization problem introduced in our paper, we extend our framework to terminal conditions without continuity or the Markovian assumption. We establish the existence and uniqueness of solutions under a smallness terminla value on the terminal conditions. We then apply this extended theory to solve a mean-field exponential utility maximization problem, which developing the classical framework of Hu et al. [Ann. Appl. Probab. 15 (2005), pp. 1691–1712] to a fully coupled quadratic mean-field setting.
Key words: Mean-field BSDEs; quadratic BSDEs; BMO martingales; Malliavin calculus; exponential utility; maximal utility.
@@ 1 Introduction
The origin of linear BSDEs can be traced back to their role as adjoint equations in stochastic control problems, first studied by Bismut (1973). The seminal paper by Pardoux and Peng (1990) established the existence and uniqueness of adapted solutions $(Y,Z)$ for
@@
under the conditions that the generator $F$ is Lipschitz continuous and the terminal condition $\xi$ is square integrable. Since then, there has been growing interest in BSDEs, with applications in stochastic control, mathematical finance, and partial differential equations. For a comprehensive presentation of the theory, see El Karoui et al. (1997), Cohen and Elliott (2015), or Zhang (2017). Numerous studies have aimed to improve existence and uniqueness results by relaxing the Lipschitz condition on the generator $F$ or the conditions on the terminal condition $\xi$.
@@
BSDEs with generators of quadratic growth in the variable $z$ have attracted significant attention. These equations arise, for example, in stochastic linear-quadratic control with random coefficients Bismut (1976), in utility optimization problems in an incomplete market Hu et al. (2005). Kobylanski (2000) provided the seminal existence and uniqueness result for real-valued quadratic BSDEs with bounded terminal values, where $F$ has quadratic growth with respect to its $Z$ component. Briand and Elie (2013) have continued to explore this area, demonstrating the existence of unique solutions to quadratic BSDEs with bounded terminal conditions using purely probabilistic methods. Investigations by Briand and Hu (2006) extend the previous results to unbounded terminal conditions. Some studies the multidimensional cases: see Hu and Tang (2016); Tang and Yang (2023); Luo (2020); Tevzadze (2008).
@@
Mean-field stochastic differential equations, also known as McKean-Vlasov equations, describe stochastic systems whose evolution is influenced by both the microscopic state and the macroscopic distribution of particles. These equations date back to Kac (1956)’s work on the Boltzmann and Vlasov equations in the 1950s. Mean-field game theory began with the groundbreaking contributions of Lasry and Lions (2007). Building on this foundation, Buckdahn et al. (2009a) introduced mean-field BSDEs, derived from approximations involving weakly interacting $N$ particle systems, and investigated them using probabilistic methods. This has led to a growing interest in mean-field BSDEs and FBSDEs. For instance, Buckdahn et al. (2009b) studied the existence and uniqueness of solutions with Lipschitz continuous generators and square integrable terminal conditions. More recently, Hao et al. (2025) explored the solvability of solutions for one-dimensional mean-field BSDEs with quadratic growth. In contrast, Tang and Yang (2023) established the existence and uniqueness of solutions for multidimensional BSDEs with generators depending on the expectation of both variables. Li et al. (2018) examined mean-field BSDEs whose coefficients depend not only on the solution $(Y,Z)$ but also on the law of $Y$. Buckdahn et al. (2026) derived a global stochastic maximum principle for mean-field forward–backward stochastic control problems with quadratic BSDE generators via BMO-based analysis.
@@
El Karoui et al. (1997) and Buckdahn et al. (2009b) state that, in a Clark–Ocone type formula, the martingale density component $Z$ of the solution pair of a BSDE with a classical globally Lipschitz generator on a Gaussian basis is the Malliavin derivative of the other component $Y$ (see Proposition 5.3). The result has been extending our understanding of quadratic BSDEs: for example, Briand and Elie (2013) provided a straightforward approach to the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus to overcome difficulties with quadratic generators, and Imkeller and Dos Reis (2010) proposed a numerical scheme for BSDEs with generators of quadratic growth.
@@
Continuous time portfolio optimization under the utility-maximization framework introduced by Merton (1969) is a foundational topic in mathematical finance. Merton analyzed the problem in a complete market where the risky asset follows the Black-Scholes dynamics and no constraints are imposed on admissible trading strategies. A major development in the early 2000s was the emergence of a connection between this class of optimization problems and the theory of BSDEs, which has since become a central approach in the field. The core idea of the BSDE methodology is to construct a stochastic process, depending on the investment strategy, that matches the investor’s utility of terminal wealth. In this context, Rouge and El Karoui (2000) used BSDEs to compute the value function and optimal strategy for exponential utility, assuming trading strategies are restricted to a convex cone. Building on this framework, Hu et al. (2005) developed a BSDE approach to utility maximization in incomplete markets and extended the methodology to include logarithmic and power utility functions, allowing for general strategy constraints within a closed set.
@@
Our methods are inspired by the works of Tevzadze (2008) and Briand and Elie (2013). The former studied the existence and uniqueness of solutions to multidimensional quadratic BSDEs using BMO martingale theory. The latter provided a straightforward approach to establishing the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus. Building on these foundations, we bridge the gap between the approaches of Cheridito and Nam (2017) and Hao et al. (2025). Specifically, Hao et al. (2025) established the existence and uniqueness by applying a contraction mapping with the Reverse Hölder Inequality (RHI) and iterating backward in time when the generator is quadratic in $Z$ component, while Cheridito and Nam (2017) used the Krasnoselskii fixed-point theorem on a small time interval, iterating backward to obtain a global solution. Our contribution extends the results of Cheridito and Nam (2017) by demonstrating in Proposition 4.9 that the function $f^{1}$ satisfies the conditions outlined in Hao et al. (2025), thereby proving the existence and uniqueness of a solution for quadratic mean-field BSDEs, which is quadratic both in $Z$ and the mean-field component $\tilde{Z}$.
Our main contributions are as follows. First, we prove existence and uniqueness for quadratic mean-field BSDEs when the generator satisfies the unified quadratic condition, that is, quadratic both in $Z$ and its mean-field component $\tilde{Z}$. In this fully coupled regime, the generator may contain genuine interactions between the control variable $Z$ and the mean-field component $\tilde{Z}$. To the best of our knowledge, this is the first paper that studies such a case in a general framework. Under this condition, we prove well-posedness for bounded terminal conditions of the form $\xi=\phi(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{n}})$, where $\phi$ is $\alpha$-Hölder continuous for $\alpha\in(0,1]$, $X$ is a solution of a forward stochastic differential equation, and $t_{i}\in[0,T]$. Second, under the separable quadratic condition (i.e. the growth of the generator with respect to $\tilde{Z}$ can be bounded by a quadratic function of $\tilde{Z}$), we obtain existence and uniqueness of the form $\xi=\phi(W_{[0,T]})$, where $\phi:C([0,T];\mathbb{R}^{d})\to\mathbb{R}$ is uniformly continuous in the sup norm. Finally, in the application part of the paper, we consider the class of generators of the same form as unified quadratic condition up to constants, and we also prove existence and uniqueness under a smallness condition on the terminal value $\xi$, without imposing additional structural regularity. Our approach builds on BMO martingale theory and Malliavin calculus to address analytical difficulties arising from the quadratic structure, and it avoids fixed-point arguments commonly used in previous studies, such as Jiang et al. (2024) and Tevzadze (2008). A key novelty throughout is the inclusion of a quadratic mean-field component $\tilde{Z}$, which substantially changes the structure compared to classical quadratic (MF)BSDEs.
@@
The rest of the paper is organized as follows. In Section 2, we present the framework, introduce the standing assumptions (including $(\mathrm{H}_{q}^{\mathrm{unif}})$ and $(\mathrm{H}_{q}^{\mathrm{sep}})$), and state our main results. Section 3 derives an almost sure upper bound for the solution pair $(Y,Z)$ of Lipschitz mean-field BSDEs, assuming the terminal condition admits a uniformly bounded Malliavin derivative, as shown in Proposition 3.1. Importantly, the resulting bound is independent of the Lipschitz constants in $z$ and $\tilde{z}$. In the main Section 4, we provide a priori estimates and establish a stability result tailored to mean-field BSDEs with quadratic growth, which will be used to pass from Lipschitz truncations to quadratic generators. In Section 5, we combine the uniform Lipschitz bounds obtained in Section 3 with the stability result and a density argument to prove existence and uniqueness under $(\mathrm{H}_{q}^{\mathrm{unif}})$ and $(\mathrm{H}_{q}^{\mathrm{sep}})$. Finally, in Section 6, we study the class of generators of the form $(\mathrm{H}_{q}^{\mathrm{LK}})$, which is equivalent to $(\mathrm{H}_{q}^{\mathrm{unif}})$ and show how the corresponding mean-field BSDE characterizes an exponential utility maximization problem under a suitable class of strategy constraints.
@@ 2.1 Preliminaries
  • $\mathbb{H}^{p}$: the set of predictable $\mathbb{R}^{d}$ valued process $Z$ satisfying $\left\|Z\right\|_{\mathbb{H}^{p}}:=\left(\mathbb{E}\left[\int_{0}^{T}\left|Z_{r}\right|^{2}\,dr\right]^{\frac{p}{2}}\right)^{\frac{1}{p}}<\infty,$
@@
    $\mathbb{E}_{\tau}\left[\int_{\tau}^{T}\left|Z_{r}\right|^{2}\,dr\right]\leq C^{2},$
    where $\mathbb{E}_{\tau}$ is the conditional expectation given $\mathcal{F}_{\tau}$. The $\left\|Z\right\|_{\mathbb{H}^{2}_{BMO}}:=\left\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\right\|_{BMO}$ is defined as the smallest non-negative constant $C$ for which the above inequality is satisfied.
@@
When a different underlying measure other than $\mathbb{P}$ is involved, we indicate it explicitly in the notation, for example by writing $\mathbb{E}$$^{\mathbb{Q}}$$\cdot$] and $\|\cdot\|_{BMO(\mathbb{Q})}$ for expectations and BMO norms taken with respect to a probability measure $\mathbb{Q}$. Here are some inequalities from BMO martingales theory, all the proofs can be found in Kazamaki (2006).
@@
For $K>0$, there exist constants $c_{1}(K),c_{2}(K)>0$, depending only on $K$, such that the following holds. Let $N$ be a one-dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$ with $\|N\|_{\mathrm{BMO}(\mathbb{P})}\leq K$, and define the equivalent measure $\tilde{\mathbb{P}}$ by
@@ 1. $(\mathrm{H}_{0})$
  2. $(\mathrm{H}_{q}^{\mathrm{unif}})$
    There exist constants $K_{z}>0$ and $K_{\tilde{z}}>0$ such that:
[…]
  3. $(\mathrm{H}_{q}^{\mathrm{sep}})$
@@
Remark 2.4.
The unified quadratic condition $(\mathrm{H}_{q}^{\mathrm{unif}})$} is the fully coupled case, in contrast with the separately quadratic $(\mathrm{H}_{q}^{\mathrm{sep}})$}. It is immediate that $(\mathrm{H}_{q}^{\mathrm{unif}})$} implies $(\mathrm{H}_{q}^{\mathrm{sep}})$}, any statement proved under $(\mathrm{H}_{q}^{\mathrm{unif}})$} also holds under $(\mathrm{H}_{q}^{\mathrm{sep}})$}, while the converse is not true. We provide an explicit example here. For simplicity, consider the one-dimensional case and define
This $g$ satisfies $(\mathrm{H}_{q}^{\mathrm{unif}})$}, but does not satisfy $(\mathrm{H}_{q}^{\mathrm{sep}})$}.
@@
The following $g$ satisfies the condition $(\mathrm{H}_{q}^{\mathrm{sep}})$}:
@@
where $l$ is locally Lipschitz in (z,z~)($\tilde{z}$) with quadratic growth in $z$ and linear growth in $\tilde{z}$, and $q$ is quadratic polynomials in $\tilde{z}$. Note that this example does not satisfy the existence and uniqueness assumption in previous literature, such as Hao et al. (2025).
Let us state our main theorems. The first main theorem assumes a Markovian structure with $(\mathrm{H}_{q}^{\mathrm{unif}})$} while the second one assumes non-Markovian structure with $(\mathrm{H}_{q}^{\mathrm{sep}})$}.
@@
Assume that there is a constant $C$ such that $b:[0,T]\times\mathbb{R}^{k}\to\mathbb{R}^{k}$ and $\sigma:[0,T]\to\mathbb{R}^{k\times n}$ are measurable functions with
  • $|b(t,x)-b(t,x^{\prime})|\leq C|x-x^{\prime}|$, $|b(t,0)|\leq C$ for every $(t,x)\in[0,T]\times\mathbb{R}^{k}$, and
  • $C^{-1}|v|^{2}\leq v^{\intercal}(\sigma\sigma^{\intercal})(t)v\leq C|v|^{2}$ for all $v\in\mathbb{R}^{k}$ and $t\in[0,T]$.
Then, let $X$ be the strong solution of
$dX_{t}=b(t,X_{t})dt+\sigma(t)dW_{t};\qquad X_{0}=x_{0}.$
Let us assume $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{unif}})$}. In addition, assume that
$\xi=\phi(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{n}}),$
for a bounded function $\phi:(\mathbb{R}^{d})^{n}\to\mathbb{R}$ which is $\alpha$-Hölder continuous, that is, there exists $K>0$ and $\alpha\in(0,1]$ such that
$|\phi(x)-\phi(x^{\prime})|\leq K\max_{i=1,2,\cdots,n}|x_{i}-x^{\prime}_{i}|^{\alpha},$
for any $x,x^{\prime}\in(\mathbb{R}^{d})^{n}$. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
@@
Note that our terminal condition is not necessarily Malliavin differentiable if $\alpha<1$.
Theorem 2.8.
Let Assumptions $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{sep}})$} hold. In addition, assume that there exists a function $\phi:C([0,T];\mathbb{R}^{d})\to\mathbb{R}$ such that $\xi=\phi(W_{[0,T]})$ and uniformly continuous in sup norm. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
Remark 2.9.
In the statement and proof of our main theorem, we may, without loss of generality, assume that
@@
Given any generator $g$ satisfying Assumptions $(\mathrm{H}_{0})$$(\mathrm{H}_{q}^{\mathrm{unif}})$}, we may introduce the shifted generator
@@
$Y_{t}=U_{t}+\int_{t}^{T}\widetilde{{\mathbb{E}}}\big[g(s,0,\tilde{Z}_{s})\big]\,ds$